# Non-Clifford and parallelizable fault-tolerant logical gates on constant and almost-constant rate homological quantum LDPC codes via higher symmetries

### v2

@@ Abstract
We study parallel fault-tolerant quantum computing for families of homological quantum low-density parity-check (LDPC) codes defined on 3-manifolds with constant or almost-constant encoding rate. We derive generic formula for a transversal $T$ gate on color codes defined on general 3-manifolds, which acts as collective non-Clifford logical CCZ gates on any triplet of logical qubits with their logical-$X$ membranes having a $\mathbb{Z}_{2}$ triple intersection at a single point. The triple intersection number is a topological invariant, which also arises in the path integral of the emergent higher symmetry operator in a topological quantum field theory (TQFT): the $\mathbb{Z}_{2}^{3}$ gauge theory. Moreover, the transversal $S$ gate of the color code corresponds to a higher-form symmetry in TQFT supported on a codimension-1 submanifold, giving rise to exponentially many addressable and parallelizable logical CZ gates. A construction of constant-depth circuits of the above logical gates via cup product cohomology operation is also presented for three copies of identical toric codes on arbitrary 3-manifolds. We have developed a generic formalism to compute the triple intersection invariants for 3-manifolds, with the structure encoded into an interaction hypergraph which determines the logical gate property and also corresponds to the hypergraph magic state that can be injected into the code. We also study the scaling of the Betti number and systoles with volume for various 3-manifolds, which translates to the encoding rate and distance. We further develop three types of LDPC codes supporting such logical gates: (1) A quasi-hyperbolic code from the product of 2D hyperbolic surface and a circle, with almost-constant rate $k/n=O(1/\log(n))$ and $O(\log(n))$ distance; (2) A homological fibre bundle code from twisting the product by an isometry of the surface based on the construction by Freedman-Meyer-Luo, with $O(1/\log^{\frac{1}{2}}(n))$ rate and $O(\log^{\frac{1}{2}}(n))$ distance; (3) A specific family of 3D hyperbolic codes: the Torelli mapping torus code, constructed from mapping tori of a pseudo-Anosov element in the Torelli subgroup, which has constant rate while the distance scaling is currently unknown. We then show a generic constant-overhead scheme for applying a parallelizable universal gate set with the aid of logical-$X$ measurements.
@@ I Introduction
Indeed, the quest of parallel schemes is quite challenging since qLDPC codes typically are composed of a single or constant number of code blocks in order for the encoding rate to be constant. This means that it can be difficult to address logical qubits individually. For instance, conventional schemes using conventional transversal logical gates tend to induce collective logic gates to many logical qubits without individual addressability of a single or constant number of logical qubits. Measurement-based schemes with individual addressability based on logical measurements have been proposed, but these schemes use additional ancilla qubits with $O(d^{2})$ overhead over a long time, where $d$ is the code distance [13]. A parallelizable scheme has recently been proposed using constant-rate concatenated codes [14]. However, this scheme uses codes with high-weight stabilizers and therefore goes beyond the scope of qLDPC codes. It remains an open question how to do individually addressable and parallelizable logical gates with qLDPC codes, without introducing additional overhead.
@@
Since the logical gate structure is completely topological, we further explore its deep origin from (3+1)D topological quantum field theory (TQFT), more specifically a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ gauge theory equivalent to three copies of 3D toric codes following the formalism developed in Ref. [29]. We then see that the logical CCZ and CZ operators correspond to emergent higher symmetries [30, 31, 32, 33, 34, 29] in TQFT, which has been a very active direction in the fields of high-energy and condensed matter physics in recent years. In a $(D+1)$-dimensional TQFT, one can think of applying a non-trivial symmetry operation on a $(D-k+1)$-dimensional submanifold by sweeping a codimension-$k$ defect along some closed trajectory in time. A special class of defects that are invertible define a higher group symmetry. Invertible codimension-$k$ defects define a $(k-1)$-form symmetry [30]. In a $(D+1)$-dimensional TQFT, this leads to a series of higher groups $K_{k}$, for $k=1,\dots,D$. This higher group structure has a correspondence with the Clifford hierarchy in quantum information in various cases [35, 36, 37, 38, 39, 29, 40, 41].
@@
In the (3+1)D case, the invertible defects are expected to define a categorical 3-group symmetry [42, 43]. The invertible codimension-$k$ defects corresponding to logical CCZ and CZ operators are so-called gauged symmetry-protected topological (SPT) defects [44, 37, 38, 39, 45, 46, 29, 40], which can be obtained from gauging a trivial phase with a SPT phase decorated on a codimension-$(k-1)$ submanifold. Specifically, the logical CCZ operator is a 00-form symmetry corresponding to sweeping a codimension-1 gauged $\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SPT defect, while the logical CZ operator is a $1$-form symmetry corresponding to sweeping a codimension-2 gauged $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SPT defect. When these symmetries are applied on a generic 3-manifold, we see that the triple intersection topological invariants naturally arise in the TQFT path integral of these symmetry operators which completely determines the logical gate structure independent of the microscopic details of the lattice models, i.e., whether they are a single 3D color code or 3 copies of 3D toric codes or some other variants. While the 0-form symmetry acts on the entire system just like the usual transversal logical gate and hence induces a collective non-Clifford gate, the 1-form (higher-form) symmetry only acts on a codimension-1 submanifold, giving rise to individually addressable and parallelizable logical Clifford gates. We note that although the idea of higher-group and higher-form symmetries in TQFT has been connected to quantum information in previous works [44, 37, 38] and has also been applied to fault-tolerant quantum computing with fractal topological codes [47], this is, as far as the authors know, the first work which finds significant advantage of applying higher symmetries in the context of quantum LDPC codes.
@@
Moreover, we use a operator-valued cochain formalism orignated from the TQFT to explicitly construct a constant-depth circuit composed of overlaping CCZ gates acting on three identical copies of toric codes defined on aribtrary triangulation on arbitrary 3-manifolds from a specific type of cohomology operation in algebraic topology: the triple cup product. We then use the operator commutation relations, the Leibniz rule of cup product, and the cocycle condition to show that such a circuit is indeed a logical gate which preserves the code space. We further show that this circuit induces a logical CCZ gate coupling a triple of logical qubits in three copies of identical toric codes when the corresponding logical-X membranes have a non-trivial triple intersection. Similarly, we derive the addressable logical CZ gate between two copies of toric codes. More elaborated study and application of this formalism on a much wider class of logical gates will discussed in the upcoming paper [48].
@@
The structure of the paper is the following: Sec. II shows the construction of color codes and surface codes on arbitrary 3-manifolds and derive the transversal T and S gates as the collective logical CCZ and parallelizable logical CZ gates respectively, as well as the connection to the triple-intersection number. One also introduces the interaction hypergraph which encodes the logical gate structure and the connection to hypergraph magic state injection. Sec. III connects the logical gates to emergent $k$-form symmetries in the topological quantum field theory and reveals the origin of the triple intersection invariant from the path integral expression using an operator-valued cochain formalism. Mathematically, this just corresponds to cohomology operation in algebraic topology. We also use this formalism to explicitly construct a constant-depth circuit composed of CCZ or CZ gates in three identical copies of toric codes and show that they lead to collective logical CCZ gates and addressable logical CZ gates. Sec. IV connects the scaling of the code parameters of the homological qLDPC codes such as the encoding rate and codes distance to the geometry of the underlying manifolds, i.e., the Betti number and systole scaling with the volume. Sec. V gives the explicit construction of the three types of homological qLDPC codes with constant or almost-constant encoding rate. Sec VI shows how to perform parallel universal fault-tolerant quantum computation with the non-Clifford and parallelizable logical gates acting on the homological qLDPC codes constructed in the previous sections. Additional logical operations including logical measurements and thickened Dehn twists are needed to achieve the universality and enhance the parallelizability. Sec.VII discusses how to back-engineer the 3-manifolds from desired logical gate structure, where the discussion focuses on the topology rather than the geometry of the 3-manifolds. The paper is concluded by Sec. VIII with the discussion of the main results and outlook of future directions.
@@ II Non-Clifford and parallelizable logical gates of color codes on 3D manifolds
II.1 Color codes and surface codes on 3D manifold
We can define a $D$-dimensional color code on some arbitrary $D$-manifold $\mathcal{M}^{D}$ [36] that is equivalent to $D$ copies of the $D$-dimensional toric-code model on the same manifold up to a constant-depth local circuit [24]. Let us review how to obtain how to obtain the color-code model from an arbitrary cellulation of $\mathcal{M}^{D}$.
@@ II.1 Color codes and surface codes on 3D manifold
Here it will be convenient to define the color code on its dual lattice with a triangulation composed of tetrahedra $\Delta$ where a tetrahedron is specified by a list of $j+1$ vertices lying on the extremities of its convex hull, i.e., $\Delta=\{v_{1},v_{2},\dots,v_{j}\}$. The lattice is $D+1$-colourable such that no two vertices of the same colour share an edge. We show how to construct such a lattice shortly, but for the present definition we will assume that such a lattice can exist. Qubits are assigned to the $D$-cells of the $D+1$-colourable lattice, and Pauli-X (Pauli-Z)-type stabilizers are associated to the 0- ($(D-2)$-) cells of the color-code lattice. Let us denote the 0-cells ($(D-2)$-cells) as $v$ ($\delta$), such that we have stabilizers
@@
where $X_{\Delta}$ ($Z_{\Delta}$) denotes a Pauli-X (Pauli-Z) operator acting on the qubit at $D$-cell $\Delta$, and we have used the subscript notation $\Delta\supset v$ ($\Delta\supset\delta$) to denote the product over all qubits $\Delta=\{u_{1},u_{2},\dots,u_{D+1}\}$ that include all the vertices of simplex $v=\{v\}$ ($\delta=\{v_{1},\dots,v_{D-1}\}$). We note that in general we can define Pauli-X and Pauli-Z stabilizers with respect to $x$- and $z$-cells, $\delta_{x}$ and $\delta_{z}$, respectively, with $x+z=D-2$ [36, 61], but for our purposes we can concentrate on $x=0$ and $z=1$ for the $D=3$ case.
We can produce a color-code lattice $\mathcal{L}_{c}$ given some arbitrary triangulation of a 3-manifold using a fattening procedure. This was shown first by Bombin and Delgado, see Appendix A of Ref. [22], but we follow the description on the dual lattice $\mathcal{L}_{c}^{*}$ due to Breuckmann and Vuillot, see Ref. [23]. We take some triangulation of a $D$ manifold, consisting of $j$-cells $\sigma_{j}$ and show we can obtain a color-code lattice as defined above. Once again, we denote the cells $\sigma_{j}$ of the $D$-dimensional manifold by a list of $D+1$ 0-cells $\sigma_{0}$. To distinguish the initial manifold that we map onto a $D+1$ colourable cellulation of the same manifold that we use to define the color code, we denote $j$-cells of the initial triangulation with the notation $\sigma_{j}$. In contrast we have denoted the $D$-cells, $(D-2)$-cells and 00-cells with symbols $\Delta$, $\delta$ and $v$, respectively, for the case of the color code lattice.
@@
where $SS\SS$ represents the stabilizer groups of decoupled ancilla qubits and $TC(\mathcal{L}_{i})$ represents the stabilizer group of the toric code defined on the shrunk lattice $\mathcal{L}_{i}$ obtained from $\mathcal{L}_{c}$ by shrinking $3$-cells of color $\text{C}_{i}$.
@@
Now we define the $i^{\text{th}}$ $\mathbb{Z}_{2}$-Betti number of $\mathcal{M}^{3}$ as $b_{i}(\mathcal{M}^{3};\mathbb{Z}_{2})$$=$$Rank(H_{i}(\mathcal{M}^{3};\mathbb{Z}_{2}))$  [^2] For conciseness, we drop $\mathbb{Z}_{2}$ in the followed texts and when we say Betti number we will always refer to the $\mathbb{Z}_{2}$ Betti number. The number of encoded logical qubits $k$ equal to both the 1st and 2nd Betti number in this case, i.e., $k=b_{1}(\mathcal{M}^{3};\mathbb{Z}_{2})=b_{2}(\mathcal{M}^{3};\mathbb{Z}_{2})$, where the second equality is a manifestation of the Poincaré duality. The code space of the 3D color code defined on manifold $\mathcal{M}^{3}$ is hence

### v3

@@ Abstract
We study parallel fault-tolerant quantum computing for families of homological quantum low-density parity-check (LDPC) codes defined on 3-manifolds with constant or almost-constant encoding rate. We derive generic formula for a transversal $T$ gate on color codes defined on general 3-manifolds, which acts as collective non-Clifford logical CCZ gates on any triplet of logical qubits with their logical-$X$ membranes having a $\mathbb{Z}_{2}$ triple intersection at a single point. The triple intersection number is a topological invariant, which also arises in the path integral of the emergent higher symmetry operator in a topological quantum field theory (TQFT): the $\mathbb{Z}_{2}^{3}$ gauge theory. Moreover, the transversal $S$ gate of the color code corresponds to a higher-form symmetry in TQFT supported on a codimension-1 submanifold, giving rise to exponentially many addressable and parallelizable logical CZ gates. A construction of constant-depth circuits of the above logical gates via cup product cohomology operation is also presented for three copies of identical toric codes on arbitrary 3-manifolds. We have developed a generic formalism to compute the triple intersection invariants for 3-manifolds, with the structure encoded into an interaction hypergraph which determines the logical gate property and also corresponds to the hypergraph magic state that can be injected into the code without distillation (‘magic state fountain’). We also study the scaling of the Betti number and systoles with volume for various 3-manifolds, which translates to the encoding rate and distance. We further develop three types of LDPC codes supporting such logical gates: (1) A quasi-hyperbolic code from the product of 2D hyperbolic surface and a circle, with almost-constant rate $k/n=O(1/\log(n))$ and $O(\log(n))$ distance; (2) A homological fibre bundle code from twisting the product by an isometry of the surface based on the construction by Freedman-Meyer-Luo, with $O(1/\log^{\frac{1}{2}}(n))$ rate and $O(\log^{\frac{1}{2}}(n))$ distance; (3) A specific family of 3D hyperbolic codes: the Torelli mapping torus code, constructed from mapping tori of a pseudo-Anosov element in the Torelli subgroup, which has constant rate while the distance scaling is currently unknown. We then show a generic constant-overhead scheme for applying a parallelizable universal gate set with the aid of logical-$X$ measurements.
@@ I Introduction
Indeed, the quest of parallel schemes is quite challenging since qLDPC codes typically are composed of a single or constant number of code blocks in order for the encoding rate to be constant. This means that it can be difficult to address logical qubits individually. For instance, conventional schemes using global or fold transversal logical gates [13, 14, 15, 16, 17, 18] tend to induce collective logic gates to many logical qubits in a code block. With the exception of topological codes that encode only a small number of logical qubits [13, 19, 14, 15, 20], few, or no examples are known for transversal gates can address individual logical qubits for codes that encode a large number of logical qubits in a single code block. Measurement-based schemes with individual addressability based on logical measurements have been proposed, but these schemes use additional ancilla qubits with $O(d^{2})$ overhead over a long time, where $d$ is the code distance [21]. A parallelizable scheme has recently been proposed using constant-rate concatenated codes [22]. However, this scheme uses codes with high-weight stabilizers and therefore goes beyond the scope of qLDPC codes. It remains an open question how to do individually addressable and parallelizable logical gates with qLDPC codes, without introducing additional overhead.
@@
Since the logical gate structure is completely topological, we further explore its deep origin from (3+1)D topological quantum field theory (TQFT), more specifically a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ gauge theory equivalent to three copies of 3D toric codes following the formalism developed in Ref. [32]. We then see that the logical CCZ and CZ operators correspond to emergent higher symmetries [33, 34, 35, 36, 37, 32] in TQFT, which has been a very active direction in the fields of high-energy and condensed matter physics in recent years. In a $(D+1)$-dimensional TQFT, one can think of applying a non-trivial symmetry operation on a (Dk+1)(D-k+1)-dimensional submanifold by sweeping a codimension-$k$ defect along some closed trajectory in time. A special class of defects that are invertible define a higher group symmetry. Invertible codimension-$k$ defects define a (k1)(k-1)-form symmetry [33]. In a $(D+1)$-dimensional TQFT, this leads to a series of higher groups $K_{k}$, for $k=1,\dots,D$. This higher group structure has a correspondence with the Clifford hierarchy in quantum information in various cases [38, 39, 40, 41, 20, 32, 42, 43].
@@
In the (3+1)D case, the invertible defects are expected to define a categorical 3-group symmetry [44, 45]. The invertible codimension-$k$ defects corresponding to logical CCZ and CZ operators [19, 20, 46, 47] are so-called gauged symmetry-protected topological (SPT) defects [48, 40, 41, 49, 32, 42], which can be obtained from gauging a trivial phase with a SPT phase decorated on a codimension-(k1)(k-1) submanifold. Specifically, the logical CCZ operator is a 0-form symmetry corresponding to sweeping a codimension-1 gauged $\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SPT defect, while the logical CZ operator is a $1$-form symmetry corresponding to sweeping a codimension-2 gauged $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SPT defect. When these symmetries are applied on a generic 3-manifold, we see that the triple intersection topological invariants naturally arise in the TQFT path integral of these symmetry operators which completely determines the logical gate structure independent of the microscopic details of the lattice models, i.e., whether they are a single 3D color code or 3 copies of 3D toric codes or some other variants. While the 0-form symmetry acts on the entire system just like the usual transversal logical gate and hence induces a collective non-Clifford gate, the 1-form (higher-form) symmetry only acts on a codimension-1 submanifold, giving rise to individually addressable and parallelizable logical Clifford gates. We note that although the idea of higher-group and higher-form symmetries in TQFT has been connected to quantum information in previous works [48, 40, 41] and has also been applied to fault-tolerant quantum computing with fractal topological codes [50], this is, as far as the authors know, the first work which finds significant advantage of applying higher symmetries in the context of high-rate quantum LDPC codes.
@@
Moreover, we use an operator-valued cochain formalism orignated from the TQFT to explicitly construct a constant-depth circuit composed of overlapping CCZ gates acting on three identical copies of toric codes defined on arbitrary triangulation on arbitrary 3-manifolds from a specific type of cohomology operation in algebraic topology: the triple cup product. We then use the operator commutation relations, the Leibniz rule of cup product, and the cocycle condition to show that such a circuit is indeed a logical gate which preserves the code space. We further show that this circuit induces a logical CCZ gate coupling a triple of logical qubits in three copies of identical toric codes when the corresponding logical-X membranes have a non-trivial triple intersection. Similarly, we derive the addressable logical CZ gate between two copies of toric codes. More elaborated study and application of this formalism on a much wider class of logical gates are discussed in the follow-up papers [51, 52].
@@
The structure of the paper is the following: Sec. II shows the construction of color codes and surface codes on arbitrary 3-manifolds and derive the transversal T and S gates as the collective logical CCZ and parallelizable logical CZ gates respectively, as well as the connection to the triple-intersection number. One also introduces the interaction hypergraph which encodes the logical gate structure and the connection to hypergraph magic state injection. Sec. III connects the logical gates to emergent $k$-form symmetries in the topological quantum field theory and reveals the origin of the triple intersection invariant from the path integral expression using an operator-valued cochain formalism. Mathematically, this just corresponds to cohomology operation in algebraic topology. We also use this formalism to explicitly construct a constant-depth circuit composed of CCZ or CZ gates in three identical copies of toric codes and show that they lead to collective logical CCZ gates and addressable logical CZ gates. Sec. IV connects the scaling of the code parameters of the homological qLDPC codes such as the encoding rate and codes distance to the geometry of the underlying manifolds, i.e., the Betti number and systole scaling with the volume. Sec. V gives the explicit construction of the three types of homological qLDPC codes with constant or almost-constant encoding rate. Sec VI shows how to perform parallel universal fault-tolerant quantum computation with the non-Clifford and parallelizable logical gates acting on the homological qLDPC codes constructed in the previous sections. Additional logical operations including logical measurements and thickened Dehn twists are needed to achieve the universality and enhance the parallelizability. The paper is concluded by Sec. VII with the discussion of the main results and outlook of future directions. Finally, we also discuss how to back-engineer the 3-manifolds from desired logical gate structure in Appendix A, where the discussion focuses on the topology rather than the geometry of the 3-manifolds.
@@ II Non-Clifford and parallelizable logical gates of color codes on 3D manifolds
In this section, we develop the theory of non-Clifford and parallelizable logical gates in the context of 3D color codes defined on 3-manifolds, while we reserve the equivalent description of three copies of 3D toric codes to Sec. III.
We first introduce the construction of color codes and toric codes on general 3-manifolds in Sec. II.1, and connect the code space to the homology group and Betti numbers. We then derive a generic formula for the collective logical CCZ gates implemented by the transversal $T$ gate and its connection to the triple-intersection invariant in the underlying 3-manifold in Sec. II.2. Next, we derive the generic formula for addressable and parallelizable logical CZ gates implemented by transversal S gates supported on a submanifold in Sec. II.3, where we have also provided the formula connecting the total number addressable logical CZ gates to the 1st Betti number in the underlying 3-manifold. Finally, in Sec. II.4, we introduce the interaction hypergraph to help understanding the structure of the logical non-Clifford gates. We further introduce the single-shot injection protocol of hypergraph magic states, which can also be represented by the interaction hypergraph.
II.1 Color codes and toric codes on 3D manifold
We can define a $D$-dimensional color code on some arbitrary $D$-manifold $\mathcal{M}^{D}$ [39] that is equivalent to $D$ copies of the $D$-dimensional toric-code model on the same manifold up to a constant-depth local circuit [13]. Let us review how to obtain the color-code model from an arbitrary cellulation of $\mathcal{M}^{D}$.
@@ II.1 Color codes and toric codes on 3D manifold
Here it will be convenient to define the color code on its dual lattice with a triangulation composed of tetrahedra $\Delta$ where a tetrahedron is specified by a list of $j+1$ vertices lying on the extremities of its convex hull, i.e., $\Delta=\{v_{1},v_{2},\dots,v_{j}\}$. The lattice is $D+1$-colourable such that no two vertices of the same colour share an edge. We show how to construct such a lattice shortly, but for the present definition we will assume that such a lattice can exist. Qubits are assigned to the $D$-cells of the $D+1$-colourable lattice, and Pauli-X (Pauli-Z)-type stabilizers are associated to the 0- ((D2)(D-2)-) cells of the color-code lattice. Let us denote the 0-cells ((D2)(D-2)-cells) as $v$ ($\delta$), such that we have stabilizers
@@
where $X_{\Delta}$ ($Z_{\Delta}$) denotes a Pauli-X (Pauli-Z) operator acting on the qubit at $D$-cell $\Delta$, and we have used the subscript notation $\Delta\supset v$ ($\Delta\supset\delta$) to denote the product over all qubits $\Delta=\{u_{1},u_{2},\dots,u_{D+1}\}$ that include all the vertices of simplex $v=\{v\}$ ($\delta=\{v_{1},\dots,v_{D-1}\}$). We note that in general we can define Pauli-X and Pauli-Z stabilizers with respect to $x$- and $z$-cells, $\delta_{x}$ and $\delta_{z}$, respectively, with x+z=D2x+z=D-2 [39, 65], but for our purposes we can concentrate on $x=0$ and $z=1$ for the $D=3$ case.
We can produce a color-code lattice $\mathcal{L}_{c}$ given some arbitrary triangulation of a 3-manifold using a fattening procedure. This was shown first by Bombin and Delgado, see Appendix A of Ref. [30], but we follow the description on the dual lattice $\mathcal{L}_{c}^{*}$ due to Breuckmann and Vuillot, see Ref. [31]. We take some triangulation of a $D$ manifold, consisting of $j$-cells $\sigma_{j}$ and show we can obtain a color-code lattice as defined above. Once again, we denote the cells $\sigma_{j}$ of the $D$-dimensional manifold by a list of $D+1$ 0-cells $\sigma_{0}$. To distinguish the initial manifold that we map onto a $D+1$ colourable cellulation of the same manifold that we use to define the color code, we denote $j$-cells of the initial triangulation with the notation $\sigma_{j}$. In contrast we have denoted the $D$-cells, (D2)(D-2)-cells and 0-cells with symbols $\Delta$, $\delta$ and $v$, respectively, for the case of the color code lattice.
@@
where $S\SS$ represents the stabilizer groups of decoupled ancilla qubits and $TC(\mathcal{L}_{i})$ represents the stabilizer group of the toric code defined on the shrunk lattice $\mathcal{L}_{i}$ obtained from $\mathcal{L}_{c}$ by shrinking $3$-cells of color $\text{C}_{i}$.
@@
Now we define the $i^{\text{th}}$ $\mathbb{Z}_{2}$-Betti number of $\mathcal{M}^{3}$ as $b_{i}(\mathcal{M}^{3};\mathbb{Z}_{2})$$Rank(H_{i}(\mathcal{M}^{3};\mathbb{Z}_{2}))$  [^2] For conciseness, we drop $\mathbb{Z}_{2}$ in the followed texts and when we say Betti number we will always refer to the $\mathbb{Z}_{2}$ Betti number. The number of encoded logical qubits $k$ equal to both the 1st and 2nd Betti number in this case, i.e., $k=b_{1}(\mathcal{M}^{3};\mathbb{Z}_{2})=b_{2}(\mathcal{M}^{3};\mathbb{Z}_{2})$, where the second equality is a manifestation of the Poincaré duality. The code space of the 3D color code defined on manifold $\mathcal{M}^{3}$ is hence